# Tap into Reality: Understanding the Impact of Interactions on Presence and Reaction Time in Mixed Reality

### v1

@@ Introduction
In Mixed Reality (MR), creating a sense of presence is essential to providing an immersive and engaging experience. Presence, the psychological sensation of being there, allows users to feel fully absorbed in a virtual environment (VE) [52]. Measuring presence is essential for improving it, and traditional approaches have primarily relied on subjective questionnaires. However, recent research has identified correlations between presence and objective measures, such as reaction time [12, 13]. These correlations have been mainly explored when presence is influenced by technical factors such as the appearance and behavior of virtual elements  [73, 59] or by conditioning-related human factors [29, 70].
@@
While these findings have provided valuable insights into the nature of presence, prior work often assumes that specific interaction mechanisms do not affect the relationship between presence and reaction time. However, literature [67, 40, 22, 68, 32, 35], has suggested that interaction design itself influences presence. In MR, interaction refers to the ways in which users engage and manipulate the VE. It encompasses the mechanisms and interfaces that allow users to perceive, navigate, and affect the virtual world. Biocca’s work [9] on interactions in VE suggests that users build a mental model of interaction patterns, and any deviation from expected behavior can affect presence. Similarly, expectations about the realism and consistency of interactions are shown to be essential in establishing a presence in MR environments [46, 31, 29]. For example, natural interactions may lead to greater presence, while unfamiliar or inconsistent interactions can disrupt presence [36, 57]. On the other hand, reaction time is typically measured as a response to a cue, often occurring as an interaction. Therefore, any approach to measuring presence should be sensitive to interactions, as it directly influences reaction time and impacts presence [54].
@@
To address this gap in the literature, we conducted an exploratory within-subjects study ($N=50$). We evaluated the relationship between presence and reaction time across two interaction mechanisms: direct (D) and symbolic (S). Direct interaction closely mimics real-world actions, like reaching out to grab or touch a virtual object. In contrast, symbolic interaction uses abstract mechanisms, such as pressing a button or performing a gesture that does not replicate real-world actions but achieves the intended effect. We examined two tasks, manipulation and selection, equally amenable to D and S, to capture a broader range of interactions. The manipulation task involves manipulating virtual objects directly, whereas the selection task demands precise coordination between user intentions and the virtual object. We followed a 2 (interaction: direct vs. symbolic) $\times$ 2 ( tasks: manipulation vs. selection) study design to create four conditions: direct manipulate (DM), symbolic manipulate (SM), direct select (DS), symbolic select (SS). We used post-experience questionnaires to measure the change in presence between conditions and systematically measured the reaction time of users in response to a visual stimulus. Based on the collected data, we aimed to answer the following research questions (RQ):
@@ 1. RQ1:
    How does the choice of interaction impact presence?
@@ 2. RQ2:
    Does the nature of the task to be performed impact the effect of interaction on presence?
@@ 1.1 Measuring Presence in MR
Our definition of MR aligns with Augmented Reality (AR) on Milgram’s reality-virtuality spectrum [39], where virtual elements merge seamlessly with the physical world. The core of this experience is “presence,” the feeling of “being there,” which relies on interactions that reflect real-world behavior [52]. Accurately measuring presence in any VE (AR/VR/MR) presents unique challenges. Traditional post-experience questionnaires [73, 54, 53, 19, 61, 48, 51, 16], often used to measure presence, capture retrospective impressions of sensory fidelity and engagement but can lack real-time insights and may suffer from memory biases [55]. Originally designed for fully immersive VR [65], questionnaires have been adapted for AR/MR [66, 71] but still primarily assess peripheral factors rather than intrinsic presence [18, 45, 65, 49, 64]. Despite limitations, they are useful across the reality-virtuality spectrum [69] if all users experience the same VE.
@@ 1.2 Presence and Interaction
Despite evidence suggesting that interaction design influences presence, research on the impact of interaction has yielded mixed results. While some studies found that high-fidelity interactions enhance immersion and enjoyment [67, 40, 22, 32, 35, 58], others reported similar outcomes between high- and low-fidelity interactions [48, 68]. Others have either focused on the sensor side of interaction [40] or explored interaction modalities [34], but none have directly addressed the impact of presence. Furthermore, many of these studies have only examined the interaction effect incidentally, treating it as a secondary research question or control measure rather than a central focus. These inconsistencies underscore the need for dedicated research to comprehensively understand the relationship between interaction and presence.
@@ 1.4 Contributions Beyond Related Work
    Task-Specific Analysis of Interaction Effects: Previous studies have not explored in detail how specific tasks interact with different interaction types to affect presence and reaction time. Our study goes beyond general findings by analyzing the impact of these specific tasks on user experience (presence) and performance behavior (reaction time). This task-specific focus provides a better understanding of how different interactions affect users depending on the nature of the task.
@@ 2.1 Theoretical Model
Reaction Time. is the duration a user takes to respond to a cue.
Presence. is the psychological sensation of “being there” within the VE [69]. Presence is about more than high-fidelity graphics; it requires a well-designed VE that replicates real-world experiences.
@@
Symbolic Interaction (§) refers to actions that do not directly replicate real-world action but instead rely on symbolic or abstract representations to achieve the desired result. For instance, using a button to select a virtual object instead of touching it. Although less intuitive, these interactions are still effective in enabling users to interact with the VE. Although D interactions are considered ideal due to their realism [7], they require more time and cost to develop. S interaction can sometimes be just as effective [44].
D is ideal for precision tasks, like assembling components by touching and aligning with natural behaviors. S is practical for tasks needing less effort, and frequent actions, like sorting items via a button, are helpful when direct interaction is not feasible. D and S are chosen for their foundational roles in MR, providing a solid base before extending research to other types. We avoid broader interactions like voice, gaze, and haptics to limit our scope and maintain clear conclusions.
Tasks. represents a specific activity or goal users aim to accomplish in VE. It is an action-oriented step or process that the user undertakes while interacting with VE. Bowman et al. [11] categorized interaction techniques into selection, manipulation, travel, and system control tasks.  [30] expanded on these by introducing canonical manipulation tasks: selection, positioning, rotation, and scaling. They highlighted the importance of using low-level tasks to evaluate interactions’ impact. The following tasks are chosen based on their prevalence and relevance in everyday virtual interactions [7, 27].
Selection tasks involve picking or highlighting objects in the VE. For instance, clicking on or tapping an object to mark it as selected or initiating an actuation. The simplicity of the task usually means it requires less effort and physical demand.
@@
Unknown. It remains unclear precisely how different types of interactions affect reaction times. The context of how interaction influences user cognitive and motor response that leads to high or low reaction time needs further exploration. The connection between specific combinations of interactions and tasks and how they impact the correlation between presence and reaction time is yet to be fully understood. This gap includes questions like: How do different tasks modify the influence of various interaction types on presence and reaction time? What patterns emerge from specific combinations of tasks and interactions that influence reaction time?
@@ 2.2 Hypothesis
  • H1: Manipulating interactions leads to a change in the presence.
    • H1.1: Participants have a higher sense of presence when experiencing direct interactions than symbolic interactions.
@@
    • H3.2: The reaction time magnitude depends on the interaction type.
@@ 2.3 Experimental Design
In the symbolic selection (SS) condition, when the color of the bubble matches the highlighted box, the participants are asked to perform the airtap interaction [1] on the highlighted box to pop the bubble, as shown in Figure 3(right). The participants point toward the box using their index finger, which creates a ray depicted as a dashed line. A donut-shaped cursor appears on the box being pointed at, signifying where the participant is aiming. Then, the participants move their index finger toward their thumb, forming a pinch gesture. As the participants transition to a pinch gesture, the dashed ray changes to a solid line, and the donut cursor shrinks to a dot, indicating that they are preparing to make a selection. Once the pinch gesture is sustained, the solid ray disappears entirely, and the participants can select the box. The selection is finalized when the pinch gesture is complete. The placement of the bubble and boxes and interaction in SS is carefully designed so that both the tasks in Select have the same physical workload and arm movement. Although the box is farther than the bubble, the intentional ray formation of the line of sight with the box compensates for the delta in the distance. We use Fitts’s law [20] to calculate the expected time of motor movement for both positions and corroborated that empirically during our pilot study (§3.5). In our design, the hypothesis assumption is that selection tasks benefit from direct touch interactions that may improve presence and are likely to reduce cognitive load. S using an anchor/hook can be efficient for rapid selections despite some cognitive disconnect. Manipulation tasks, like grabbing and rotating, are more intuitive with D and could increase the presence. S interactions may require learning but can become efficient once mastered.
@@ 2.4 Interaction Implementation
Interactions in this study are managed through the HoloLens’ built-in hand tracking (accuracy of Hololens2 hand tracking is empirically tested by [26, 63]), allowing participants to use their hand’s gestures to move the box (Manipulate) and pop bubble (Select), as they would in the physical world. Pointing and clicking gestures are excluded from the D conditions to maintain realism. Each gesture step flows into the next with tactile feedback (cursor or ray or color change) to the participants so that gestures are interpreted correctly by the system. Following HoloLens 2’s near-interaction guidelines, primary objects (boxes and bubbles) are placed 35-50 cm from the participant’s abdomen for ease of access and are ergonomically positioned to prevent muscle fatigue [21]. To reduce accidental selections, participants are taught specific hand gestures consistent with the standard HoloLens 2 hand gestures interaction paradigm [3]. Lastly, the box is placed on the table surface in Manipulate conditions to compensate for the haptic feedback.
@@ 3.1 Participants
We conducted an a priori power analysis to determine the appropriate sample size for our within-subject four conditions (DM, SM, DS, SS) study for medium size effect ($0.25$[14] with four repeated measures, an $\alpha=0.05$, a power of $0.95$, and nonsphericity correction ($\epsilon=1$), the analysis recommended a minimum of 39 participants. We recruited 50 participants to account for attrition (33 male-identifying, 17 female-identifying) with a mean age of 21.86 years (ST = 3.39). All participants provided written informed consent and received \$15 for their participation. All the participants had normal or corrected normal vision with contact lenses or glasses. The demographic distribution and media usage of participants can be seen in Table 1. The institution’s ethics committee approved the study.
@@ 3.2 Material
We selected Hololens 2 [2] due to its direct view of the real world, a self-contained device with a holographic processing unit with a Qualcomm Snapdragon 850 CPU, featuring eye, spatial, and hand-tracking capabilities and dual displays with a $1440\times 936$ pixels resolution with 110°FoV. The VEs were developed using the Unity game engine and C# programming language with the Unity XR Plug-in Management package, optimized for the Universal Windows Platform. Custom scripts were added to facilitate hand tracking and reaction time recording.
@@ 3.3 Experimental Task
In Section §2.3, we explain the experimental conditions and designs. Here, we will summarize the physical tasks (move, touch, grab, point), virtual object placements, and how they are ordered in the study session. The study features four conditions to evaluate two interactions across two tasks (as shown in Figure Tap into Reality: Understanding the Impact of Interactions on Presence and Reaction Time in Mixed Reality). For two Select task conditions, we designed a color-matching activity involving a bubble and four boxes. Participants stand as the scene is rendered before them, with bubbles positioned at their eye level and adjusted for height. Every 3s (seconds), the bubble changes color, and every 9s, the box color changes. Participants must pop the bubble when its color matches the box’s, completing 20 bubble-popping trials in 3 minutes. The delay and duration were empirically adjusted during the pilot study (§3.5). In the D, participants pop bubbles by directly touching them, while in the S, they use an air-tap gesture to interact with the boxes (§2.3).
@@
For two Manipulate task conditions, we designed a box moving activity using a single virtual box. Participants are asked to sit in a chair positioned in front of a table and adjust the chair so they can comfortably reach the box. Every 5 seconds, the box turns green, signaling the participant to move the box to the side. Starting from a neutral position, they move the box to the right in one trial, then to the left in the following trial, continuing this alternating pattern for 3 minutes across 35 trials. To prevent participants from moving the box too far, white tape marks were pasted on the table 8 inches apart to indicate the target areas. The distance of movement was empirically determined during the pilot study. In the D interaction, participants use both hands to move the box with a real-world mimicking grab gesture directly. In contrast, in the S interaction, they use a pinch gesture to manipulate hooks on the box (§2.3).
@@
In all conditions, we measure the reaction time as an interval between the time the color change occurred (cue) and the actual action taken by the participant (interaction). The feedback is explicit in all conditions: successful actions are marked by either the bubble popping or the box changing position, depending on the specific condition. To build a comprehensive response time profile, we concurrently recorded timestamps of all events, including color changes, touch, movement, and transitions, in a separate thread. Additionally, we logged the distance each participant moved the box, identified which box was moved, and verified if it matched the highlighted box. This comparison determined the accuracy: a perfect match indicated the correct color pairing. While memory retention is not a focus of this study, we log this data to exclude response times for incorrect color matches, ensuring comparability across all participants and conditions.
3.4 Measurements
We focus on presence and reaction time as the primary variables in this study to investigate how they relate to different interaction mechanisms and tasks, as outlined in our hypotheses (§2.2). We measured presence using two post-experience questionnaires: the Witmer and Singer Presence Questionnaire (PQ)[73] and the Igroup Presence Questionnaire (IPQ)[54, 53]. The PQ evaluates factors such as the possibility to act and examine, realism, self-evaluation, and interface quality. The IPQ measures three factors: the sense of physically being in the VE, realism, and involvement (crucial for evaluating interactions). The presence scores are derived from 33 items (14 IPQ, 19 PQ) on a 7-point scale with no modifications. Additionally, we included a prompt question, ”How connected do you feel to this virtual world?” on a 7-point scale, created by [13], to measure presence within the scene between trials in all four conditions and to align with reaction time. Presence is the dependent variable in H1, H1.1, H2, H2.1, where interaction type (independent variable) and task type (moderating variable in H2) are expected to affect the participants’ sense of presence.
Our software recorded reaction time on the HoloLens2 in milliseconds. We collected 20 reaction time measurements for each participant for the Select conditions and 35 for the Manipulate conditions. Reaction time is the dependent variable in H3, where we examine its relationship with presence and how it might vary based on interaction type (independent variable in H3.1) and the magnitude of this effect (examined in H3.2).

### v2

@@ Introduction
In Mixed Reality (MR), creating a sense of presence is essential to providing an immersive and engaging experience. Presence, the psychological sensation of being there, allows users to feel fully absorbed in a virtual environment (VE) [55]. Presence can take various forms, including spatial presence (the feeling of physically existing in the VE) [17], social presence (the perception of interaction with others) [46], and self-presence (the sense of embodiment and identity within the VE) [52, 57]. This study measures spatial presence, aligning with the tasks designed to assess user interactions in VE. Measuring presence is essential for improving it, and traditional approaches have primarily relied on subjective questionnaires. However, recent research has identified correlations between presence and objective measures, such as reaction time [13, 14]. These correlations have been mainly explored when presence is influenced by technical factors such as the appearance and behavior of virtual elements  [77, 62] or by conditioning-related human factors [32, 74].
@@
While these findings have provided valuable insights into the nature of presence, prior work often assumes that specific interaction mechanisms do not affect the relationship between presence and reaction time. However, literature [71, 42, 25, 72, 35, 38], has suggested that interaction design itself influences presence. In MR, interaction refers to the ways in which users engage and manipulate the VE. It encompasses the mechanisms and interfaces that allow users to perceive, navigate, and affect the virtual world. Biocca’s work [10] on interactions in VE suggests that users build a mental model of interaction patterns, and any deviation from expected behavior can affect presence. Similarly, expectations about the realism and consistency of interactions are shown to be essential in establishing a presence in MR environments [48, 34, 32]. For example, natural interactions may lead to greater presence, while unfamiliar or inconsistent interactions can disrupt presence [39, 60]. On the other hand, reaction time is typically measured as a response to a cue, often occurring as an interaction, and has been shown to vary based on scene fidelity [13] and user conditioning [14]. Therefore, any approach to measuring presence should be sensitive to interactions to provide accurate experience assessments, as it directly influences reaction time and impacts presence [57].
@@
To address this gap in the literature, we conducted an exploratory within-subjects study ($N=50$). We evaluated the relationship between presence and reaction time across two interaction mechanisms: direct (D) and symbolic (S). Direct interaction closely mimics real-world actions, such as reaching out to grab or touch a virtual object. Symbolic interaction, on the other hand, often employs abstract mechanisms such as pressing a button or performing a gesture that does not replicate real-world actions but achieves the intended effect. Still, symbolic gestures may also involve degrees of familiarity, depending on user conditioning and task goals [16]. We examined two tasks, manipulation and selection, equally amenable to D and S, to capture a broader range of interactions and to assess the joint effects of interaction type and task type. The manipulation task involves manipulating virtual objects directly, whereas the selection task demands precise coordination between user intentions and the virtual object. We followed a 2 (interaction: direct vs. symbolic) $\times$ 2 ( tasks: manipulation vs. selection) study design to create four conditions: direct manipulate (DM), symbolic manipulate (SM), direct select (DS), symbolic select (SS). We used post-experience questionnaires to measure the change in presence between conditions and systematically measured the reaction time of users in response to a visual stimulus. The goal was to understand how these choices influenced presence and reaction time. Based on the collected data, we aimed to answer the following research questions (RQ):
@@ 1. RQ1:
    To what extent does interaction type influence presence in MR?
@@ 2. RQ2:
    How does the task type modulate the impact of interaction type on presence?
@@ 1.1 Measuring Presence in MR
Our definition of MR aligns with Augmented Reality (AR) on Milgram’s reality-virtuality spectrum [41], where virtual elements merge seamlessly with the physical world. The core of this experience is “presence," the feeling of “being there," which relies on interactions that reflect real-world behavior [55]. To refine the analysis, this work focuses on spatial presence, which is primarily influenced by the user’s perception of their location and interaction within the VE. Self and social presence are outside the scope of this study but remain relevant for multi-user and embodiment-focused MR experiences [21]. Accurately measuring presence in any VE (AR/VR/MR) presents unique challenges. Traditional post-experience questionnaires [77, 57, 56, 22, 64, 50, 54, 18], often used to measure presence, capture retrospective impressions of sensory fidelity and engagement but can lack real-time insights and may suffer from memory biases [58]. Originally designed for fully immersive VR [69], questionnaires have been adapted for AR/MR [70, 75] but still primarily assess peripheral factors rather than intrinsic presence [21, 47, 69, 51, 68]. Despite limitations, they are useful across the reality-virtuality spectrum [73] if all users experience the same VE.
@@ 1.2 Presence and Interaction
Despite evidence suggesting that interaction design influences presence, research on the impact of interaction has yielded mixed results. While some studies found that high-fidelity interactions enhance immersion and enjoyment [71, 42, 25, 35, 38, 61], others reported similar outcomes between high- and low-fidelity interactions [50, 72]. Others have either focused on the sensor side of interaction [42] or explored interaction modalities [37], but none have directly addressed the impact of presence. Many of these studies have only examined the interaction effect incidentally, treating it as a secondary research question or control measure rather than a central focus. These inconsistencies underscore the need for dedicated research to comprehensively understand the relationship between interaction and presence.
@@ 1.4 Contributions Beyond Related Work
    Task-Specific Analysis of Interaction Effects: Previous studies have not explored in detail how specific tasks interact with different interaction types to affect presence and reaction time. Our study complements existing findings by analyzing the impact of these particular tasks, task-specific interactions, and their joint effects on user experience (presence) and performance behavior (reaction time). This task-specific focus provides a better understanding of how different interactions affect users depending on the nature of the task.
@@ 2.1 Theoretical Model
Presence. is the psychological sensation of “being there" within the VE [73]. Presence is about more than high-fidelity graphics; it requires a well-designed VE that replicates real-world experiences.
@@
  • Symbolic Interaction (S) refers to actions that do not directly replicate real-world action but rely on symbolic or abstract representations to achieve the desired result. For instance, using a button to select a virtual object instead of touching it. Although less intuitive, these interactions are still effective in enabling users to interact with the VE. Although D interactions are preferred due to their realism [8], they can be time-consuming and costly, requiring higher computational overhead. In contrast, S interactions (raycasts or button presses) offer practical advantages in scenarios requiring fast, repetitive inputs, reduced physical effort, and interactions beyond arm’s reach [9].
D is considered beneficial for precision-oriented tasks, such as assembly tasks that align with familiar motor behaviors [12]. S is practical for tasks needing less effort, and frequent actions, like sorting items via a button, are helpful when direct interaction is not feasible [8]. D and S are chosen for their foundational roles in MR, providing a solid base before extending research to other types. We limit our scope by avoiding broader interactions like voice, gaze, and haptics.
Tasks. represents a specific activity or goal users aim to accomplish in VE. It is an action-oriented step or process that the user undertakes while interacting with VE. Bowman et al. [12] categorized interaction techniques into selection, manipulation, travel, and system control tasks. LaViola et al. [33] expanded on these by introducing canonical manipulation tasks: selection, positioning, rotation, and scaling. They highlighted the importance of using low-level tasks to evaluate interactions’ impact. The following tasks are chosen based on their prevalence and relevance in everyday virtual interactions [8, 30].
  • Selection tasks involve choosing or highlighting objects in the VE. For instance, clicking on or tapping an object to mark it as selected or initiating an actuation. The simplicity of the task usually means it requires less effort and physical demand.
@@
Unknown. It remains unclear precisely how different interaction types impact reaction times. Further exploration is required to determine how much interaction influences cognitive and motor responses, resulting in faster or slower reaction times. The connection between specific combinations of interaction types and task types and how they impact the correlation between presence and reaction time is yet to be fully understood.
@@ 2.2 Hypothesis
  • H1: Manipulating interactions (direct and symbolic) leads to a change in the presence.
    • H1.1: Participants experience a higher sense of presence during direct interactions than symbolic interactions.
@@
    • H3.2: The extent of the change in reaction time depends on the interaction type.
@@ 2.3 Experimental Design
In the symbolic selection (SS) condition, when the color of the bubble matches the highlighted box, the participants are asked to perform the airtap interaction [1] on the highlighted box to pop the bubble, as shown in Figure 3(right). The participants point toward the box using their index finger, which creates a ray depicted as a dashed line. A donut-shaped cursor appears on the box being pointed at, signifying where the participant is aiming. Then, the participants move their index finger toward their thumb, forming a pinch gesture. As the participants transition to a pinch gesture, the dashed ray changes to a solid line, and the donut cursor shrinks to a dot, indicating that they are preparing to make a selection. Once the pinch gesture is sustained, the solid ray disappears entirely, and the participants can select the box. The selection is finalized when the pinch gesture is complete. The placement of the bubble and boxes and interaction in SS is carefully designed so that both the tasks in Select have the same physical workload and arm movement. Although the box is farther than the bubble, the intentional ray formation of the line of sight with the box compensates for the delta in the distance. We use Fitts’s law [23] to calculate the expected time of motor movement for both positions and corroborated that empirically during our pilot study (§3.5). In our design, the hypothesis assumption is that selection tasks benefit from direct touch interactions that may improve presence and are likely to reduce cognitive load. S using an anchor/hook can be efficient for rapid selections despite some cognitive disconnect. Manipulation tasks, like grabbing and rotating, are more intuitive with D and could increase the presence. S interactions may require learning but can become efficient once mastered. In the selection task, participants selected objects using either a button press (S) or by reaching out to touch the target (D). These differing interaction modes reflect the distinct mechanics of symbolic and direct interaction in MR. However, we recognize that the S could have been implemented using a more consistent ray-cast targeting method rather than a button press. The choice of a button press was informed by its common usage in MR tasks involving symbolic actions but introduces a limitation due to the significant difference in selection targets between the two interaction types.
@@ 2.4 Interaction Implementation
Interactions in this study are managed through the HoloLens’ built-in hand tracking (accuracy of Hololens2 hand tracking is empirically tested by [29, 67]), allowing participants to use their hand’s gestures to move the box (Manipulate) and pop bubble (Select), as they would in the physical world. Pointing and clicking gestures are excluded from the D conditions to maintain realism. Each gesture step flows into the next with tactile feedback (cursor or ray or color change) to the participants so that gestures are interpreted correctly by the system. Following HoloLens 2’s near-interaction guidelines, primary objects (boxes and bubbles) are placed 35-50 cm from the participant’s abdomen for ease of access and are ergonomically positioned to prevent muscle fatigue [24]. To reduce accidental selections, participants are taught specific hand gestures consistent with the standard HoloLens 2 hand gestures interaction paradigm [4]. Lastly, in Manipulate conditions, the box was placed on the table to provide haptic feedback.
@@ 3.1 Participants
We conducted an a priori power analysis to determine the appropriate sample size for our within-subject four conditions (DM, SM, DS, SS) study for medium size effect ($0.25$[15] with four repeated measures, an $\alpha=0.05$, a power of $0.95$, and nonsphericity correction ($\epsilon=1$), the analysis recommended a minimum of 39 participants. We recruited 50 participants to account for attrition (33 male-identifying, 17 female-identifying) with a mean age of 21.86 years (ST = 3.39). All participants were recruited through the college campus. All participants provided written informed consent and received \$15 for their participation. All the participants had normal or corrected vision with contact lenses or glasses and the absence of motion sickness history. The demographic distribution and media usage of participants can be seen in Table 1. The study was approved by the UMASS Amherst IRB, Human Research Protection Office.
@@ 3.2 Material
We selected Hololens 2 [3] due to its direct view of the real world, a self-contained device with a holographic processing unit with a Qualcomm Snapdragon 850 CPU, featuring eye, spatial, and hand-tracking capabilities and dual displays with a $1440\times 936$ pixels resolution with 110°FoV. The VEs were developed using the Unity game engine and C# programming language with the Unity XR Plug-ins, optimized for the Universal Windows Platform. Custom scripts were added to facilitate hand tracking and reaction time recording.
@@ 3.3 Experimental Task
In Section §2.3, we explain the experimental conditions and designs. Here, we will summarize the physical tasks (move, touch, grab, point), virtual object placements, and how they are ordered in the study session. The study features four conditions to evaluate two interactions across two tasks (as shown in Figure Tap into Reality: Understanding the Impact of Interactions on Presence and Reaction Time in Mixed Reality). We designed a color-matching activity involving a bubble and four boxes for two Select task conditions. Participants stand as the scene is rendered before them, with bubbles positioned at their eye level and adjusted for height. Every 3s (seconds), the bubble changes color, and every 9s, the box color changes. Participants must pop the bubble when its color matches the box’s, complete 20 bubble-popping trials, and set the total phase duration to 3 minutes. In the D, participants pop bubbles by directly touching them, while in the S, they use an air-tap gesture to interact with the boxes (§2.3). Here, the phase length is determined by time rather than the exact number of trials, ensuring a fixed session length across participants while maintaining consistency in the task structure.
@@
For two Manipulate task conditions, we designed a box moving activity using a single virtual box. Participants are asked to sit in a chair positioned in front of a table and adjust the chair so they can comfortably reach the box. Every 5 seconds, the box turns green, signaling the participant to move the box to the side. Starting from a neutral position, they move the box to the right in one trial and then to the left in the following trial, continuing this alternating pattern. Participants perform 35 trials of moving the box within a phase that also lasts 3 minutes. The phase duration is again set by time, with the number of trials serving as a reference point for participants to maintain pace. To prevent participants from moving the box too far, white tape marks were pasted on the table 8 inches apart to indicate the target areas. The distance of movement was empirically determined during the pilot study. In the D interaction, participants use both hands to move the box with a real-world mimicking grab gesture directly. In contrast, in the S interaction, they use a pinch gesture to manipulate hooks on the box (§2.3). The durations for both tasks were determined during pilot testing to balance participant engagement and avoid fatigue (§3.5).
@@
In all conditions, we measure the reaction time as an interval between the time the color change occurred (cue) and the actual action taken by the participant (interaction). Reaction time was chosen as an objective performance metric based on prior findings linking faster responses to increased presence [13, 14]. This metric complements subjective presence scores, providing a real-time proxy for cognitive engagement. The feedback is explicit in all conditions: successful actions are marked by either the bubble popping or the box changing position, depending on the specific condition. To build a comprehensive response time profile, we concurrently recorded timestamps of all events, including color changes, touch, movement, and transitions, in a separate thread. Additionally, we logged the distance each participant moved the box, identified which box was moved, and verified if it matched the highlighted box. This comparison determined the accuracy: a perfect match indicated the correct color pairing. While memory retention is not a focus of this study, we log this data to exclude response times for incorrect color matches, ensuring comparability across all participants and conditions.
@@
3.4 Measurements
We focus on presence and reaction time as the primary variables in this study to investigate how they relate to different interaction mechanisms and tasks, as outlined in our hypotheses (§2.2). We measured presence using two post-experience questionnaires: the Witmer and Singer Presence Questionnaire (PQ)[77] and the Igroup Presence Questionnaire (IPQ)[57, 56]. The PQ evaluates factors such as the possibility to act and examine, realism, self-evaluation, and interface quality. The IPQ measures three factors: the sense of physically being in the VE, realism, and involvement (crucial for evaluating interactions). The presence scores are derived from 33 items (14 IPQ, 19 PQ) on a 7-point scale with no modifications. Additionally, we included a prompt question, "How connected do you feel to this virtual world?" on a 7-point scale, created by [14], to measure presence within the scene between trials in all four conditions and to align with reaction time. Presence is the dependent variable in H1, H1.1, H2, H2.1, where interaction type (independent variable) and task type (moderating variable in H2) are expected to affect the participants’ sense of presence.
Our software recorded reaction time on the HoloLens2 in milliseconds. We collected 20 reaction time measurements for each participant for the Select conditions and 35 for the Manipulate conditions. Reaction time is the dependent variable in H3, where we examine its relationship with presence and how it varies based on interaction type (independent variable in H3.1) and the magnitude of this effect (examined in H3.2). As the task phases were time-based rather than trial-completion-based, the analysis accounted for any variability in the number of completed trials when calculating average reaction times.